OpenStudy (anonymous):

1/3m+1/2m+7=5/6m+5+2

OpenStudy (anonymous):

\[\frac{ 1 }{ 3m } + \frac{ 1 }{ 2m } +7 = \frac{ 5 }{ 6m } + 5 +2\] look like that?

OpenStudy (anonymous):

I gt 1 2/3 + 14 n yes

OpenStudy (anonymous):

i got 0=0

OpenStudy (anonymous):

Hw

OpenStudy (anonymous):

man i wish i could show u my paper haha

OpenStudy (anonymous):

I see now thank u !

OpenStudy (anonymous):

i made \[\frac{ 1 }{ 3m } + \frac{ 1 }{ 2m } into \frac{ 2 }{ 6m } + \frac{ 3 }{ 6m }\]